# Dialogue Systems for Emotional Support via Value Reinforcement

### v1

@@ Abstract
Emotional support dialogue systems aim to reduce help-seekers’ distress and help them overcome challenges. While human values—core beliefs that shape an individual’s priorities—are increasingly emphasized in contemporary psychological therapy for their role in fostering internal transformation and long-term emotional well-being, their integration into emotional support systems remains underexplored. To bridge this gap, we present a value-driven method for training emotional support dialogue systems designed to reinforce positive values in seekers. Our model learns to identify which values to reinforce at each turn and how to do so, by leveraging online support conversations from Reddit. The model demonstrated superior performance in emotional support capabilities, outperforming various baselines. Notably, it more effectively explored and elicited values from seekers. Expert assessments by therapists highlighted two key strengths of our model: its ability to validate users’ challenges and its effectiveness in emphasizing positive aspects of their situations—both crucial elements of value reinforcement. Our work validates the effectiveness of value reinforcement for emotional support systems and establishes a foundation for future research. [^1] This paper is currently under review. All source code and data will be made publicly available upon its publication.
@@
Juhee Kim1 Chunghu Mok2 Jisun Lee2 Hyang Sook Kim2 Yohan Jo1 Graduate School of Data Science1, Department of Psychology2 Seoul National University {inbz1244, chunghumok, jisune87, hyangkim, yohan.jo}@snu.ac.kr
Figure 1: Comparison of value-independent and value-reinforced approaches in emotional support, based on examples sourced from Reddit.
@@ 1 Introduction
Emotional support aims to help individuals (seekers) in addressing everyday emotional difficulties, such as relationship conflicts and workplace stress, by offering reassurance, acceptance, and encouragement (Atoum and Al-Shoboul, 2018; Burleson, 2003). Recent advancements in large language models have accelerated the development of dialogue systems designed to provide emotional support (supporters) (Deng et al., 2024; Zhang et al., 2023; Chen et al., 2023). Although many studies emphasize reinforcing positive emotions in seekers, this approach has limitations. Emotional changes alone may not adequately capture deeper intrinsic transformations within the seeker, potentially reducing the long-term impact of emotional support.
@@
To overcome these limitations, we propose an emotional support approach grounded in value reinforcement. Human values, which represent core beliefs and guiding principles, help individuals determine what is important and meaningful in life (Searle, 2003), such as self-direction, benevolence, tradition, etc. Given their deep connection to life purpose and personal identity, values play a central role in modern psychological interventions, such as Acceptance and Commitment Therapy (ACT) (Plumb et al., 2009; Hayes and Pierson, 2005) and Values Affirmation Interventions (Miyake et al., 2010; Jordt et al., 2017).
Incorporating values into emotional support is expected to offer several benefits, including a deeper understanding of the seeker’s transformation and the promotion of long-term emotional well-being. Assessing how positive values are reflected in a seeker’s utterances provides insights into their acceptance of support and willingness to change, as shown in Figure 1. Furthermore, the ACT approach highlights that engaging with personal values enhances clarity, direction, and motivation, empowering individuals to navigate challenges with purpose (Blackledge and Hayes, 2001). This supports the ultimate goal of achieving a healthy life—not merely feeling good but living well—thereby fostering long-lasting emotional well-being. The importance of values in emotional support is further demonstrated through the widely used emotional support dataset, ESConv (Liu et al., 2021). Our analysis reveals that positive values are more prominently expressed in the high-effectiveness group of seekers (i.e., high reduction in negative emotions) (see Figure 2 and Section 3 for details).
@@
In this paper, we present a framework for training a supporter model through simulations with a seeker simulator. To enhance the supporter’s ability to reinforce the seeker’s values, we introduce two components trained on Reddit data: (1) a target value detector that identifies the values to promote at each turn, and (2) a reference generator that generates a supporter response to reinforce these values. By integrating their outputs, the supporter model aims to maximize the reward of value promotion reflected in the seeker’s responses. The training involves two phases: supervised fine-tuning, which distills the simulation capability of GPT-4o-mini into a smaller model, and direct policy optimization (Rafailov et al., 2023), which enhances the model’s value reinforcement effectiveness.
@@
  • We propose an effective two-phase approach, featuring a target value detector and a reference generator, both trained on real-world data from Reddit.
@@ 2.2 Dialogue Systems for Emotional Support
Recent advances in language models have broadened their use in emotional support dialogue systems. To enhance supporter models, researchers have explored various approaches. One method uses large language models (LLMs) to generate diverse conversations for supporter model training (Zheng et al., 2024; Liu et al., 2023; Qiu et al., 2024). Other studies predict seekers’ future states to refine support model training (Zhou et al., 2023; Cheng et al., 2022; Shin et al., 2020). Recent efforts also leverage multi-turn simulations with user simulators to capture nuanced reactions. For instance, Deng et al. (2024) used seeker simulators to predict future responses and train policy planners guiding supporter models. However, most studies often overlook the role of human values. While our study builds on simulation-based training, its main contribution lies in integrating values into emotional support, emphasizing their critical role in improving system effectiveness.
@@ 3.2 Exploring the Impact of Values on Emotional Support Effectiveness
To motivate our research, we conducted an analysis to examine the role of values in emotional support by analyzing the ESConv dataset (Liu et al., 2021), which contains multi-turn emotional support conversations in English among crowdworkers. We analyze whether the reinforcement of a seeker’s values positively influences the effectiveness of emotional support.
@@
In ESConv, seekers rated the intensity of their negative emotions before and after the conversation on a scale from 1 (lowest) to 5 (highest). In our analysis, dialogues with an initial intensity of 5 are divided into two groups: high effectiveness (final intensity of 1–2) and low effectiveness (final intensity of 3–4). We then analyze positive value expressions in the seekers’ final four turns using automated classifiers (Liu et al., 2024; Schroter et al., 2023). Detailed experimental procedures are described in Appendix A.
Figure 2: Comparison of positive value expressions between high and low effectiveness groups, along with the average predicted probabilities per turn for values showing significant differences.
@@ Results.
In the last four turns, the high effectiveness group exhibited a significantly higher average number of positive values (7.9) compared to the low effectiveness group (6.5). Figure 2 highlights values that were pronounced in the high effectiveness group. Table 34 provides examples of seekers’ utterances that illustrate these values.
@@ 4 Emotional Support Dataset from Reddit
Our goal is to use this data to train a model that identifies the values to reinforce at each turn (target value detector) and a model that produces supporter utterances to effectively promote the target values (reference generator). So, we labeled the data with sentiment strength and expressed values at both the post and comment levels using models developed by Liu et al. (2024) and Schroter et al. (2023). Values expressed in a positive comment by the OP can be considered successful target values at that time, while the preceding comment can be regarded as an effective supporter utterance that promotes those values. The dataset contains over 20,000 samples, with details provided in the appendix B.
@@ 5 Method
The overall framework, illustrated in Figure 3, consists of three core components: (1) the target value detector, which identifies values to reinforce in the seeker at each turn; (2) the reference generator which produces utterances to effectively promote these values from the seeker; (3) the supporter model which determines strategies and generates responses based on the identified target values and the reference responses. To ensure the reproducibility of our work, we will release our source code, trained models, and simulation data publicly.
@@ 5.1 Target Value Detector
We train the target value detector using the high-quality emotional support conversations from Reddit (Section 4). Given a dialogue history $o_{1},c_{1},o_{2},c_{2},...,c_{t-1},o_{t}$, where $o_{i}$ and $c_{i}$ represent the $i$th utterances by the OP (seeker) and a commenter (supporter), respectively, the target value detector predicts which values to target in $c_{t}$. The ground-truth values $v_{t+1}$ are the top-3 values observed in $o_{t+1}$, based on their probabilities from the value detection model (Schroter et al., 2023).
@@ 5.3 Supporter Model
The supporter model is the primary model that interacts with the seeker, generating responses that align with target values. It processes three key inputs: the dialogue history, the target values identified by the target value detector at each turn, and a reference response generated by the reference generator. At each turn, the model generates reasoning across four aspects (Figure 3): (1) identifying the seeker’s issues and current state, (2) analyzing the key content of the reference response, (3) determining whether to incorporate the reference response, and (4) selecting the optimal emotional support strategy (Appendix C) and generating the subsequent response. The reason for step (3) is that, while the Reddit data offers valuable information across diverse emotional support scenarios, its distribution may not align perfectly with everyday conversations. Thus, the model selectively incorporates reference responses during reasoning.
@@
SFT requires large-scale emotional support conversations grounded in value reinforcement. To obtain such data, we opt to use dialogue simulation with a seeker simulator (Section 5.4). We use GPT-4o-mini for both the supporter and seeker simulators to generate data for training a smaller model [^3] In our pilot experiment, zero-shot Llama-3-8B-Instruct was found to be unsuitable as a supporter simulator due to issues like repetitive responses and biases in reference usage.. The simulators engage in interactions by iteratively producing an utterance based on the ongoing dialogue history as a prompt and appending it to the history prompt. The detailed prompts are provided in Table 18 and Table 28.
During simulations, GPT avoids using reference responses in approximately 90% of cases. To prevent models fine-tuned on this data from inheriting the same bias, we simulate additional responses (called “alternative responses”) at each supporter turn. Specifically, if GPT initially used the reference response, we simulate an alternative response without the reference response, and vice versa. This approach enables fine-tuned models to explore diverse dialogue flows, mitigating GPT’s biases.
The simulated dialogues are employed to fine-tune Llama-3-8B-Instruct, with dataset sizes outlined in Table 1. Detailed training methods are provided in Appendix D.3.
@@
where $N_{t+k}(v_{t+1})$ is the frequency of the values targeted at turn $t$ appearing in the seeker’s utterance at turn $t+k$, $h$ is the look-ahead horizon (the number of future steps considered), and $\gamma$ is a discount factor balancing immediate and future rewards. Response pairs are added to the DPO dataset only if their reward difference exceeds a threshold $T_{\text{diff}}$.
For the dialogues underlying the preference data above, we conduct additional dialogue simulations between the SFT supporter model and the seeker simulator. This is because the SFT model has an enhanced ability to generate and explore diverse dialogue flows. Details of the methods are provided in Appendix D.3. Table 1 summarizes the total dataset sizes, while hyperparameter details are presented in Table 19.
@@ 5.4 Seeker Simulator
To simulate various scenarios, we generated personas using GPT-4o and GPT-4o-mini, defining attributes such as problem type, emotions, and situations (Figure 3). These attributes were informed by prior studies (Liu et al., 2021; Zhao et al., 2024). The process resulted in 2,036 unique personas: 1,796 for training, 120 for development, and 120 for testing. Details on persona generation and examples can be found in Appendix E.1.
The seeker simulator is based on GPT-4o-mini. It outperformed alternative models in generating responses that align with human-like content, emotional tone, and value alignment. Experimental results and the simulator prompt are provided in Appendices E.2 and E.3, respectively.
@@
6 Experiments
6.1 Evaluation Methods
We evaluate the emotional support capabilities of various models through simulated conversations between each supporter model and the seeker simulator, using the 120 predefined seeker personas for testing. A conversation is considered complete if the seeker simulator generates “[END]” or if the seeker’s emotion score, as calculated by EmoLlama-Chat-7B (Liu et al., 2024), reaches 0.6 or higher and is accompanied by gratitude-related expressions (e.g., “thank you”). To account for practical conversation lengths, interactions are limited to a maximum of 20 turns, based on the average 15-turn length observed in the ESConv dataset. Only conversations concluding within this limit are included in the evaluation.
6.2 Evaluation Metrics
We conduct evaluations focusing on three key aspects: ES-Skills, ES-Intensity, and ES-Value.
ES-Skills evaluates a supporter’s emotional support capabilities across three components, drawing upon previous studies (Zheng et al., 2024; Zhao et al., 2024; Cheng et al., 2023; Deng et al., 2024; Cheng et al., 2022; Liu et al., 2021): (1) emotional support skills, which include Identification, Comforting, Suggestions, Experience, and Informativeness; (2) general conversation skills, covering Consistency, Role-Adherence, Expression, and Humanness; and (3) an Overall score. Each criterion is rated on a five-point scale using GPT-4o-mini. Detailed metric descriptions are in Appendix F.1.
ES-Intensity measures the intensity of a seeker’s negative emotions after a conversation. Scores are assigned on a five-point scale, with lower scores indicating minimal negative emotions. We developed a predictive model using GPT-4o-mini based on ratings provided by human seekers in ESConv. The model demonstrates a correlation of 0.345 with the actual ratings. A detailed explanation of the model and its performance is provided in Appendix F.2.
ES-Value assesses value reinforcement from two perspectives: the seeker’s experience of value exploration and reinforcement within conversations, and the supporter’s contribution to this process. We conduct pairwise comparisons between models using GPT-4o-mini as a judge. The reason is that, when assessing conversations individually on a 1–5 scale, GPT tends to award scores of 4 and 5 to most conversations, making it difficult to discern performance differences among models. Details are outlined in Appendix F.3.
To validate our GPT-based evaluation for ES-Skills and ES-Value, we calculated correlations with ratings from licensed therapists. All criteria showed positive correlations (0.198–0.778), most of which were statistically significant. Further details are provided in Appendix H.

### v2

@@ Abstract
Emotional support dialogue systems aim to reduce help-seekers’ distress and help them overcome challenges. While human values—core beliefs that shape an individual’s priorities—are increasingly emphasized in contemporary psychological therapy for their role in fostering internal transformation and long-term emotional well-being, their integration into emotional support systems remains underexplored. To bridge this gap, we present a value-driven method for training emotional support dialogue systems designed to reinforce positive values in seekers. Notably, our model identifies which values to reinforce at each turn and how to do so, by leveraging online support conversations from Reddit. We evaluate the method across support skills, seekers’ emotional intensity, and value reinforcement. Our method consistently outperforms various baselines, effectively exploring and eliciting values from seekers. Additionally, leveraging crowd knowledge from Reddit significantly enhances its effectiveness. Therapists highlighted its ability to validate seekers’ challenges and emphasize positive aspects of their situations—both crucial elements of value reinforcement. Our work, being the first to integrate value reinforcement into emotional support systems, demonstrates its promise and establishes a foundation for future research. [^1] All source code and data will be made publicly available upon its publication.
@@
Juhee Kim1 Chunghu Mok2 Jisun Lee2 Hyang Sook Kim2 Yohan Jo1thanks: Corresponding author 1Graduate School of Data Science, 2Department of Psychology Seoul National University {inbz1244, chunghumok, jisune87, hyangkim, yohan.jo}@snu.ac.kr
Figure 1: An example dialogue based on our method. The supporter model receives the target values to reinforce from the seeker and a reference response for guidance. It then selects an appropriate emotional support strategy and generates a response for the next turn.
@@ 1 Introduction
Emotional support aims to help individuals (seekers) in addressing everyday emotional difficulties, such as relationship conflicts and workplace stress, by offering reassurance, acceptance, and encouragement (Atoum and Al-Shoboul, 2018; Burleson, 2003). Recent advancements in large language models have accelerated the development of dialogue systems designed to provide emotional support (supporters) (Deng et al., 2024; Zhang et al., 2023; Chen et al., 2023). Many models have focused on reinforcing positive emotions in seekers. However, emotional changes alone may not adequately capture deeper intrinsic transformations within the seeker, potentially reducing the long-term impact of emotional support (Blackledge and Hayes, 2001). For instance, a seeker’s perfunctory “Thank you”, used as a conversational pleasantry, receives a higher positivity score (0.758) from a sentiment classifier than the response shown in Figure 1 (0.583) [^2] We employed EmoLlama-Chat-7B (Liu et al., 2024), which demonstrates superior performance in this task.. The low positivity score of the latter response may be attributed to phrases like “I know it’s not going to be easy”, which could be perceived as negative. However, the latter response demonstrates a stronger commitment and willingness to change, highlighting the importance of evaluating support effectiveness beyond emotion alone.
@@
To overcome these limitations, we propose an emotional support approach grounded in value reinforcement. Human values, which represent core beliefs and guiding principles, help individuals determine what is important and meaningful in life (Searle, 2003), such as self-direction, benevolence, tradition, etc. Given their deep connection to life purpose and personal identity, values play a central role in modern psychological interventions, such as Acceptance and Commitment Therapy (ACT) (Plumb et al., 2009; Hayes and Pierson, 2005) and Values Affirmation Interventions (Miyake et al., 2010; Jordt et al., 2017). These techniques aim to help seekers commit to goals aligned with their values, fostering intrinsic and long-term transformation. This supports the ultimate goal of achieving a healthy life—not merely feeling good but living well. The importance of values in emotional support is further demonstrated by the widely used emotional support dataset, ESConv (Liu et al., 2021). Our analysis reveals that positive values are more prominently expressed in the high-effectiveness group of seekers (i.e., high reduction in negative emotions) (see Section 3 for details).
@@
In this paper, we present a framework for training a supporter model through simulations with a seeker simulator. To enhance the supporter’s ability to reinforce the seeker’s values, we introduce two key components trained on Reddit data: (1) a target value detector that identifies the values to promote at each turn, and (2) a reference generator that generates a supporter response to reinforce these values. By integrating their outputs, the supporter model aims to maximize the reward of value promotion reflected in the seeker’s responses. Figure 1 illustrates our approach applied to an example dialogue, which reinforces the seeker’s values along with their acceptance of support and willingness to change. The training involves two phases: supervised fine-tuning, which distills the simulation capability of GPT-4o-mini into a smaller model, and direct policy optimization (Rafailov et al., 2023), which enhances the model’s value reinforcement effectiveness.
@@
  • We propose an effective two-phase approach, featuring a target value detector and a reference generator, both trained on real-world knowledge from Reddit.
@@ 2.2 Dialogue Systems for Emotional Support
To enhance supporter models, researchers have explored various approaches. One method uses large language models to generate diverse conversations for supporter model training (Zheng et al., 2024; Liu et al., 2023; Qiu et al., 2024). Other studies predict seekers’ future states to refine supporter model training (Zhou et al., 2023; Cheng et al., 2022; Shin et al., 2020). Recent efforts also leverage multi-turn simulations with seeker simulators to predict future responses (Deng et al., 2024). However, most studies often overlook the role of human values. While our study builds on simulation-based training, its main contribution lies in integrating values into emotional support, emphasizing their critical role in improving system effectiveness.
@@ 3.2 Exploring the Impact of Values on Emotional Support Effectiveness
To motivate our research, we conducted an analysis to examine the role of values in emotional support by analyzing the ESConv dataset (Liu et al., 2021), which contains multi-turn emotional support conversations in English among crowdworkers. We analyze whether reinforcing a seeker’s values positively influences the effectiveness of emotional support.
@@
In ESConv, seekers rated the intensity of their negative emotions before and after the conversation on a scale from 1 (lowest) to 5 (highest). In our analysis, dialogues with an initial intensity of 5 are divided into two groups: high effectiveness (final intensity of 1–2) and low effectiveness (final intensity of 3–4) [^4] There were no cases in the ESConv where the final emotional intensity remained at 5.. We then analyze positive value expressions in the seekers’ final four turns using automated classifiers (Liu et al., 2024; Schroter et al., 2023). This focus on the final four turns accounts for differences in turn length across groups and captures changes resulting from the emotional support conversation. Detailed experimental procedures are described in Appendix A.
Figure 2: Average number of value expressions in the last four turns in ESConv for high and low effectiveness groups.
@@ Results.
According to the analysis, the high effectiveness group exhibited a significantly higher average number of positive values (7.9) than the low effectiveness group (6.5) in the last four turns. Figure 2 highlights values that were pronounced in the high effectiveness group. Table 39 provides examples of seekers’ utterances that illustrate these values.
@@ 4 Emotional Support Dataset from Reddit
Our goal is to use this data to train a model that identifies the values to reinforce at each turn (target value detector) and a model that produces supporter utterances to effectively promote the target values (reference generator). For this purpose, we labeled the data with sentiment strength and expressed values at both the post and comment levels using models developed by Liu et al. (2024) and Schroter et al. (2023). Values expressed in a positive comment by the OP can be considered successful target values at that time, while the preceding comment from a commenter can be regarded as an effective supporter utterance that promotes those values. The dataset contains over 20,000 samples, with details on the classification models and the generated dataset provided in Appendix B.
@@ 5 Method
The overall framework, illustrated in Figure 3, consists of three core components: (1) target value detector identifies values to reinforce at each turn; (2) reference generator produces utterances to effectively promote these values from the seeker; (3) supporter model determines strategies and generates responses based on the identified target values and the reference responses. To ensure the reproducibility of our work, we will release our source code, trained models, and simulation data publicly.
@@ 5.1 Target Value Detector
We train the target value detector using the emotional support conversations from Reddit (Section 4). Given a dialogue history ($o_{1},c_{1},o_{2},c_{2},...,c_{t-1},o_{t}$), where $o_{i}$ and $c_{i}$ represent the $i$th utterances by the OP (seeker) and a commenter (supporter), respectively, the target value detector predicts which values to target in $c_{t}$. The ground-truth values $v_{t+1}$ are the top-3 values observed in $o_{t+1}$, based on their probabilities from the value detection model (Schroter et al., 2023).
@@ 5.3 Supporter Model
The supporter model is the primary model that interacts with the seeker, generating responses that align with target values. It processes three key inputs: the dialogue history, the target values identified by the target value detector at each turn, and a reference response generated by the reference generator. At each turn, the model generates a response using the reasoning process across the following four aspects (Figure 3): (1) identifying the seeker’s issues and current state, (2) analyzing the key content of the reference response, (3) determining whether to incorporate the reference response into the final output, and (4) selecting the optimal emotional support strategy (Appendix C) and generating the final response. The entire prompt is in Table 22. In step (3), the model generates either “Yes”, along with an explanation of how the reference will be incorporated, or “No”, with justification if the reference is deemed unsuitable. This selective incorporation is necessary because, while Reddit data offers valuable information across diverse emotional support scenarios, its distribution may not always align with everyday conversations. We compare our method against the direct use of Reddit-based reference responses in section 6.4.2.
@@
SFT requires large-scale emotional support conversations grounded in value reinforcement. To obtain such data, we opt to use dialogue simulation with a seeker simulator (Section 5.4). We use GPT-4o-mini for both the supporter and seeker simulators to generate data for training a smaller model [^5] In our pilot experiment, zero-shot Llama-3-8B-Instruct was found to be unsuitable as a supporter simulator due to issues like repetitive responses and biases in reference usage.. The simulators engage in interactions by iteratively producing an utterance based on the ongoing dialogue history as a prompt and appending it to the history prompt.
During simulations, GPT avoids using reference responses in approximately 90% of cases. To prevent models fine-tuned on this data from inheriting the same bias, we simulate additional responses (called “alternative responses”) at each supporter turn. Specifically, if GPT initially used the reference response, we simulate an alternative response without the reference response, and vice versa.
The simulated dialogues are employed to fine-tune Llama-3-8B-Instruct, with dataset sizes outlined in Table 1.
@@
where $N_{t+k}(v_{t+1})$ is the frequency of the values targeted at turn $t$ appearing in the seeker’s utterance at turn $t+k$, $h$ is the look-ahead horizon (the number of future steps considered), and $\gamma$ is a discount factor balancing immediate and future rewards. A response pair is added to the DPO dataset only if their reward difference exceeds a threshold $T_{\text{diff}}$.
To prepare the dialogues underlying the preference data above, we conduct additional dialogue simulations between the SFT supporter model and the seeker simulator. This is because the SFT model has an enhanced ability to generate and explore diverse dialogue flows. Table 1 summarizes the total dataset sizes, while hyperparameter details are presented in Table 23. Details of the methods are provided in Appendix D.3.
Figure 4: Win ratios in human evaluation comparing the naturalness of responses from the seeker simulator and human seekers on the ESConv (Liu et al., 2021) and AnnoMI (Wu et al., 2022, 2023) datasets.
@@ 5.4 Seeker Simulator
The seeker simulator generates seeker utterances based on the provided persona and dialogue history. To simulate various scenarios, we generated personas using GPT-4o and GPT-4o-mini, defining attributes such as problem type, emotions, and situations (Figure 3), informed by prior studies (Liu et al., 2021; Zhao et al., 2024). The process resulted in 2,036 unique personas: 1,796 for training, 120 for development, and 120 for testing.
The seeker simulator is based on GPT-4o-mini, with its detailed design and validity provided in Appendix  E. To summarize, we extensively verified the quality of the seeker simulator using human and automated evaluations. Human evaluators judged our seeker simulator to be as natural as, or more natural than, human seekers (Figure 4). Additionally, it produced utterances that more closely resemble human seekers’ responses in content, emotional tone, and value alignment, compared to other simulators.
6 Experiments
6.1 Evaluation Methods
We evaluate various supporter models through conversations with the seeker simulator, using the 120 held-out seeker personas for testing. A conversation is considered complete if the seeker simulator generates “[END]” or if the seeker’s emotion score, as calculated by EmoLlama-Chat-7B (Liu et al., 2024), reaches 0.6 or higher with gratitude expressions (e.g., “thank you”). Interactions are limited to a maximum of 20 turns, based on the average conversation length of 15 turns observed in the ESConv dataset. Only conversations concluding within this limit are included in the evaluation.
6.2 Evaluation Metrics
We conduct evaluations focusing on three key aspects: ES-Skills, ES-Intensity, and ES-Value. A detailed explanation of the metrics is in Appendix F.
ES-Skills evaluates a supporter’s capabilities across three components, based on prior studies (Zheng et al., 2024; Zhao et al., 2024; Cheng et al., 2023; Deng et al., 2024; Cheng et al., 2022; Liu et al., 2021): (1) emotional support skills, including Identification, Comforting, Suggestions, Experience, and Informativeness; (2) general conversation skills, covering Consistency, Role-Adherence, Expression, and Humanness; and (3) an Overall. Each criterion is rated on a five-point scale using GPT-4o-mini.
ES-Intensity measures the intensity of a seeker’s negative emotions after a conversation. Scores are assigned on a five-point scale, with lower scores indicating minimal negative emotions. We developed a predictive model using GPT-4o-mini based on ratings provided by human seekers in ESConv. The model demonstrates a correlation of 0.345 with the actual ratings.
ES-Value assesses value reinforcement from two perspectives: the seeker’s experience of value exploration and reinforcement within conversations, and the supporter’s contribution to this process. We conduct pairwise comparisons between models using GPT-4o-mini as a judge. The reason is that, when assessing conversations individually on a 1–5 scale, GPT tends to award scores of 4 and 5 to most conversations, making it difficult to discern performance differences among models.
To validate our GPT-based evaluation for ES-Skills and ES-Value, we calculated correlations with ratings from licensed therapists. All criteria showed positive correlations (0.198–0.778), most of which were statistically significant (Appendix K).